what is the answer -x+[x]+5
Found 2 solutions by lynnlo, JBarnum:
Answer by lynnlo(4176)
Answer by JBarnum(2146)

What are u trying to solve? What x is or if u just add -x and positive x which leaves 5
But if x needs to have a value is all numbers greater than or equal to 0 to get the answer 5 but all numbers less than 0 will come to different answers when added
-0+|0|+5=5
-5+|5|+5=5
-10+|10|+5=5
But
-(-5)+|-5|+5=15
-(-10)+|-10|+5=25
Hope this helps